OpenStudy (idealist10):

Solve the initial-value problem y'+2xy=x^2, y(0)=3 and leave the answer in a form involving a definite integral.

OpenStudy (idealist10):

@SithsAndGiggles @hartnn

OpenStudy (idealist10):

I have the work: the integrating factor u=e^x^2 (e^x^2)y'+2xy(e^x^2)=(x^2)(e^x^2) Now what do I do? How do I integrate (x^2)(e^x^2)? I did integration by parts but it only complicated the problem more.

OpenStudy (anonymous):

\[\begin{align*}e^{x^2}y'+2xe^{x^2}y&=x^2e^{x^2}\\ \frac{d}{dx}\left[e^{x^2}y\right]&=x^2e^{x^2}\\ e^{x^2}y&=\int x^2e^{x^2}~dx\\ y&=e^{-x^2}\int x^2e^{x^2}~dx\end{align*}\] The integral doesn't have an elementary result, so you can leave it as is. The final answer will involve a definite integral form, though, to account for the initial value.

OpenStudy (anonymous):

Here's some info about how to set up the integral: http://math.mit.edu/suppnotes/suppnotes03/d.pdf

OpenStudy (idealist10):

But y(0)=3, when I plug 0 into the y equation, the integral of 0 is 0, right?

OpenStudy (anonymous):

For an IVP with initial value \(y(x_0)=y_0\), the solution would have this form: \[\large e^{x^2}y=y_0+\int_{x_0}^xf(t)~dt\] so in this case, since \(y(0)=3\), you would use \[\large e^{x^2}y=3+\int_0^xt^2e^{t^2}~dt=3+\int_0^xt^2e^{t^2}~dt\] I'm actually not entirely sure about the \(\large e^{x^2}\) where it is... Let's see what happens when you check to see if this solution works. Differentiate both sides (apply the FTC for the integral): \[\large e^{x^2}y'+2xe^{x^2}y=0+x^2e^{x^2}\] which does indeed simplify to the original equation once you divide out the \(e^{x^2}\) factors.

OpenStudy (anonymous):

So to summarize, the answer would be \[\large y=3e^{-x^2}+e^{-x^2}\int_0^xt^2e^{t^2}~dt\]

OpenStudy (idealist10):

Thank you so much!

OpenStudy (anonymous):

You're welcome!